# VisCRA: A Visual Chain Reasoning Attack for Jailbreaking Multimodal Large Language Models

### v1

@@ 3.1 Vulnerability of MLRMs
Prior research in LRMs has identified a fundamental trade-off between reasoning depth and safety alignment, showing that enhanced reasoning capabilities can inadvertently weaken a model’s resistance to harmful prompts (Jiang et al., 2025). In this work, we extend this line of inquiry to MLRMs by systematically investigating how enhanced visual reasoning affects their safety.
@@ 3.2 Attack Amplification via Visual CoT
Building on the observation from Section 3.1 that stronger visual reasoning increases MLLMs vulnerability, we hypothesize that explicitly eliciting visual Chain-of-Thought (CoT) reasoning could further amplify attack success. To test this, we design a tailored visual CoT prompt to encourage step-by-step reasoning over image content (See Appendix A.1 for the prompt template.). Empirical results confirm our hypothesis: integrating
@@
visual CoT with HADES adversarial instructions significantly boosts jailbreak success rates (as illustrated in Figure 1, the increase from ’HADES’ to ’HADES+CoT’ bars for each model), highlighting the power of guided visual reasoning in bypassing safety mechanisms. However, this approach also reveals an important failure mode. While detailed image descriptions can aid reasoning, over-describing harmful visual content too early in the reasoning process can generate an excess of toxic tokens, inadvertently triggering built-in safety filters. This results in the model rejecting the harmful prompt before execution, as illustrated in Figure 2.

### v2

@@ 3.1 Vulnerability of MLRMs
Prior research in LRMs has identified a fundamental trade-off between reasoning and safety, showing that enhanced reasoning capabilities can weaken a model’s resistance to harmful prompts (Jiang et al., 2025). In this work, we extend this to MLRMs by systematically investigating how enhanced visual reasoning affects their safety.
@@ 3.2 Attack Amplification via Visual CoT
Building on the observation from Section 3.1 that stronger visual reasoning increases MLLMs vulnerability, we hypothesize that explicitly eliciting visual Chain-of-Thought (CoT) reasoning could further amplify attack success. To test this, we design a tailored visual CoT prompt to encourage step-by-step reasoning over image content
@@
(See Appendix A.1 for the prompt template.). Empirical results confirm our hypothesis: integrating visual CoT with HADES adversarial instructions significantly boosts jailbreak success rates (as illustrated in Figure 1, the increase from ’HADES’ to ’HADES+CoT’ bars for each model), highlighting the power of guided visual reasoning in bypassing safety mechanisms. However, this approach also reveals an important failure mode. While detailed image descriptions can aid reasoning, over-describing harmful visual content too early in the reasoning process can generate an excess of toxic tokens, inadvertently triggering built-in safety filters. This results in the model rejecting the harmful prompt before execution, as illustrated in Figure 2.